# Recent Progress in Flavor Model Building

### v1

@@ Abstract
quarks and leptons, flavor puzzles, beyond the SM, flavor physics
@@ 1 INTRODUCTION
Within the landscape of quantum field theories, the Standard Model (SM) of particle physics is a prominent example of a chiral, spontaneously broken non-abelian gauge theory, wherein anomaly cancellation imposes stringent constraints on the charge assignments of matter fields within a single generation. However, nature often surprises us with unexpected complexity, exemplified by the existence of three generations of particles—an apparent redundancy that has no obvious reason. This replication of fermion gauge representations is referred to as flavor. Understanding the origin of flavor remains one of the most compelling and unresolved questions in particle physics.
@@ 2 FLAVOR PUZZLES
In this section, we define and discuss the flavor puzzles of the SM and NP separately.
@@ 2.1 The Standard Model Flavor Puzzle
where $q_{p},u_{p},d_{p},\ell_{p}$, and $e_{p}$ are the five gauge representations of the chiral fermions, with $p,r=1,2,3$ as flavor indices. The core of the flavor puzzle lies in the following question: Why is there a hierarchical structure in the input parameters $Y_{f}^{pr}$ when they all enter the theory in the same way, coupling to a single Higgs field? According to the singular value decomposition theorem, $Y_{f}=L_{f}\hat{Y}_{f}R_{f}^{\dagger}$, where $\hat{Y}_{f}$ is a diagonal matrix with real and positive entries, and $L_{f}$ and $R_{f}$ are unitary matrices. The masses are determined by the singular values $\hat{M}_{f}=\hat{Y}_{f}\langle H\rangle$, while the quark mixing is given by the CKM matrix $V=L^{\dagger}_{u}L_{d}$. The experimentally determined values of the fermion masses and the CKM matrix elements, shown in Figure 1, reveal a highly hierarchical structure in the Yukawa matrices.
Figure 1: Top: Masses of Standard Model particles in GeV$/c^{2}$, including uncertainties. Values for the Higgs boson, gauge bosons, quarks, and leptons are taken from the PDG [4]. Neutrino masses, several orders of magnitude lighter, are not shown. Two neutrino mass differences have been measured [5], and cosmological data constrain the sum of neutrino masses to $\sum m_{\nu}\lesssim 0.1$ eV [6]. The lightest neutrino could potentially be massless. Bottom: Absolute values of CKM matrix elements with uncertainties, derived from direct measurements as compiled by the PDG [4].
@@
Alternatively to the neutrino mass origin outlined above, the lepton number could remain unbroken, making neutrinos Dirac particles. In this scenario, their masses would arise from renormalizable Yukawa interactions, similar to the charged fermions in Eq. (1). The small neutrino masses would then result from extremely tiny Yukawa couplings, further complicating the flavor puzzle.
Flavor model building attempts to explain the non-generic flavor structure of the SM fermions described above. Although this review focuses on the quark sector, the lack of flavor hierarchies in the neutrino sector offers essential insight into the problem.
@@ 2.2 The New Physics Flavor Puzzle
One practical implication of the SM flavor puzzle, even without knowing its solution, is the existence of approximate accidental symmetries within the SM. The smallness of Yukawa couplings and the CKM alignment leads to approximate flavor and CP conservation, resulting in highly suppressed FCNCs and EDMs. The resulting selection rules, exemplified by the GIM mechanism [11], have been instrumental in designing precision tests of the SM that are highly sensitive to new physics, which typically breaks these approximate symmetries. So far, these tests have yielded negative results. However, the null findings have profound consequences for the flavor structure of new physics if it exists near the TeV scale—a situation known as the new physics flavor puzzle. That is to say, new flavor parameters of a TeV-scale extension of the SM are also not generic.
@@
Consider a flavor-anarchic scenario in which all dimensionless flavor parameters are of order one, assuming no particular hierarchy or structure in the flavor sector of the new physics. For a tree-level matching, this predicts $c_{\mathcal{O}}\sim 1$. Under this assumption, many of the dimension-6 operators contribute significantly to processes involving neutral meson oscillations ($\Delta F=2$), charged lepton flavor violation (cLFV), EDMs, or strangeness-changing decays ($\Delta S=1$). The experimental absence of deviations from SM predictions in such processes imposes stringent constraints on the scale $\Lambda$, often pushing it to values as high as six orders of magnitude above the TeV scale. For a concrete illustration, Figure 1 of Ref. [13] presents bounds on a selected set of operators in the Warsaw basis, explicitly showing the flavor indices in the down-quark charged-lepton mass basis.
@@
The main motivation for expecting new physics at the TeV scale is the Higgs hierarchy problem, which asks why the Higgs boson mass is so much lighter than the Planck scale despite quantum corrections that should drive it higher. Symmetry-based solutions to this problem—such as supersymmetry or composite Higgs models—imply that new physics is expected not far above the EW scale to stabilize the Higgs mass. These theories often introduce new particles and interactions with order-one couplings to the Higgs field. Consequently, the flavor structure of such new physics cannot be generic; it must be hierarchical and (or) aligned to avoid excessive violation of the SM’s approximate flavor and CP symmetries. This raises the possibility that the NP and the SM flavor structure might share a common origin.
@@ 3.1.1 Symmetry-Based Models
The Froggatt-Nielsen (FN) mechanism is a seminal example utilizing an abelian horizontal symmetry, typically a $U(1)$ symmetry, to explain flavor hierarchies [14, 15, 16]. In this framework, the SM fermions carry generation-dependent charges under the additional $U(1)$ symmetry. A scalar field, known as the FN (or flavon) field ($\Phi$), acquires a vacuum expectation value (vev) $\langle\Phi\rangle$, thus spontaneously breaking the flavor symmetry. The vev $\langle\Phi\rangle$ is assumed to be smaller than the cutoff scale $\Lambda$ of the effective theory. Yukawa couplings arise from higher-dimensional operators suppressed by powers of the small parameter $\varepsilon=\langle\Phi\rangle/\Lambda\ll 1$. By assigning appropriate $U(1)$ charges, the FN mechanism can qualitatively reproduce the observed mass hierarchies and mixing angles from a single expansion parameter $\varepsilon$. For instance, heavier fermions are assigned lower charges, resulting in less suppression in their Yukawa couplings. The most straightforward UV completion of the higher dimensional operators are chains of vector-like fermions that communicate the symmetry breaking to the SM fermions. The FN setup contains a sufficient number of free $O(1)$ parameters to reproduce the observed fermion masses precisely. If the $U(1)$ symmetry is gauged, the models contain a $Z^{\prime}$ gauge boson with a mass of order $\langle\Phi\rangle$. If the $U(1)$ is global, one instead expects a very light pseudo-Nambu-Goldstone boson. These particles generically have flavor-changing couplings and can be effectively searched for with flavor probes. Flavor models commonly feature a decoupling limit: by sending both $\langle\Phi\rangle$ and $\Lambda$ to high energies while keeping $\varepsilon=\langle\Phi\rangle/\Lambda$ constant, these models can address the SM flavor puzzle without introducing excessive FCNCs at low energies. Current experimental constraints impose a lower bound on the scale of flavor dynamics, while future experiments could potentially reveal evidence for low-scale flavor models.
@@
Non-abelian flavor symmetries, such as $SU(3)$ or $SU(2)$, organize the fermion generations into non-trivial group representations. By assigning fermions to multiplets, these symmetries constrain the structure of the Yukawa couplings. The observed flavor structure typically arises through sequential and hierarchical symmetry breaking. Non-abelian symmetries often lead to more restrictive textures than FN models. However, challenges in these models include the complexity of the symmetry-breaking sector to reproduce realistic parameters. Models based on $SU(3)$ flavor symmetry have been proposed to generate realistic mass hierarchies; see e.g. [17, 18, 19, 20]. However, the sizable Yukawa coupling of the top quark ($y_{t}\sim 1$) breaks the $SU(3)$ symmetry, making $SU(2)$ (or better $U(2)$) groups more attractive [21, 22, 23, 24, 25, 26, 27]. In this construction, the first two families form a $U(2)$ doublet, while the third family is a singlet. In the symmetric limit, only the third generation acquires mass, providing a decent starting point. Symmetry breaking typically occurs in two stages: first, the $U(2)\equiv SU(2)\times U(1)$ symmetry is broken to an intermediate $U(1)$, generating a small mass for the second generation; subsequently, the $U(1)$ symmetry is broken, yielding an even smaller mass for the first generation. For recent developments, see section 4.1.
@@
Radiative models propose that lighter SM fermion masses arise from loop corrections rather than at the tree level [28]. In these scenarios, mass hierarchies arise from the suppression associated with loop factors, which intriguingly coincide with the two orders of magnitude separating the masses of adjacent generations. The absence of tree-level Yukawa couplings for the light generations can, for example, be imposed by using flavor symmetries. The breaking of a flavor symmetry is then communicated from a separate sector of the model to the SM fermions at the loop level. Additional fields and interactions are introduced to facilitate the necessary loop diagrams. In contrast to FN models, their masses can be at the same scale as flavor breaking, as the role of $\varepsilon$ is effectively replaced by a loop factor. The wide range of possible quantum numbers and interactions for the new fields gives rise to a plethora of models, both supersymmetric [29, 30, 31, 32] and non-supersymmetric [33, 34, 35, 36].
@@ 3.1.2 Geometry-Based Models
A concrete example in which this mechanism can be at work is the Randall-Sundrum (RS) model [38] that introduces a warped extra dimension with a non-factorizable geometry, where the metric depends exponentially on the extra-dimensional coordinate. The extra dimension is bounded by an infrared (IR) brane and an ultraviolet (UV) brane. The warping leads to exponential factors that naturally generate hierarchies. The Higgs naturalness problem is addressed if the Higgs resides on the IR brane. Fermions that propagate in the bulk have profiles that interpolate between the UV and IR branes and can be localized at different positions along the extra dimension, resulting in hierarchical Yukawa couplings with the Higgs. Heavier fermions, such as the top quark, are more localized towards the IR brane, resulting in larger Yukawa couplings [39, 40, 41].
The flavor-dependent, extra-dimensional profiles of the SM fermions imply a rich flavor phenomenology [42, 43]. In generic setups, the constraints from meson mixing are very stringent and push the new physics scale to $\mathcal{O}(10)$ TeV. Extra dimensional models can be protected from large FCNCs by using bulk and brane flavor symmetries [44].
Partial compositeness proposes that SM fermions are mixtures of elementary and composite states arising from a strongly coupled sector not far above the electroweak scale. Models of partial compositeness were originally introduced to tame FCNCs in technicolor models [45]. The light fermions are mostly elementary, while the heavy top quark is mostly composite. Per the AdS/CFT correspondence, models of partial compositeness map onto models with warped extra dimensions, with the amount of compositeness given by the localization of the fermions along the extra dimension.

### v2

@@ Abstract
quarks, leptons, flavor puzzles, beyond the SM, flavor physics
@@ 1 INTRODUCTION
Within the landscape of quantum field theories, the Standard Model (SM) of particle physics is a prominent example of a chiral, spontaneously broken non-abelian gauge theory, wherein anomaly cancellation imposes stringent constraints on the charge assignments of matter fields within a single generation. However, nature often surprises us with unexpected complexity, exemplified by the existence of three generations of particles—an apparent redundancy that has no obvious reason. This replication of fermion gauge representations is referred to as flavor. Understanding the origin of flavor remains a longstanding and unresolved question in particle physics.
@@ 2 FLAVOR PUZZLES
In this section, we define and discuss the SM and NP flavor puzzles separately.
@@ 2.1 The Standard Model Flavor Puzzle
where $q_{p},u_{r},d_{r},\ell_{p}$, and $e_{r}$ are the five gauge representations of the chiral fermions, with $p,r=1,2,3$ as flavor indices and $\widetilde{H}=i\sigma_{2}H^{*}$. The core of the flavor puzzle lies in the following question: Why is there a hierarchical structure in the input parameters $Y_{f}^{pr}$ when they all enter the theory in the same way, coupling to a single Higgs field? According to the singular value decomposition theorem, $Y_{f}$ equals $L_{f}\hat{Y}_{f}R_{f}^{\dagger}$, where $\hat{Y}_{f}$ is a diagonal matrix with real and positive entries, and $L_{f}$ and $R_{f}$ are unitary matrices. The masses are determined by the singular values $\hat{M}_{f}=\hat{Y}_{f}\langle H\rangle$, while the quark mixing is given by the CKM matrix $V=L^{\dagger}_{u}L_{d}$. The experimentally determined values of the fermion masses and the CKM matrix elements, shown in Figure 1, reveal a highly hierarchical structure in the Yukawa matrices.
Figure 1: Top: Masses of Standard Model particles in units of GeV$/c^{2}$, including uncertainties. Values for the Higgs boson, gauge bosons, quarks, and leptons are taken from the PDG [4]. Neutrino masses, several orders of magnitude lighter, are not shown. Two neutrino mass differences have been measured [5], and cosmological data constrain the sum of neutrino masses to $\sum m_{\nu}\lesssim 0.1$ eV [6]. The lightest neutrino could potentially be massless. Bottom: Absolute values of CKM matrix elements with uncertainties, derived from direct measurements as compiled by the PDG [4]. Figure adapted from Reference [7], with permission from the author.
@@
As an alternative to the neutrino mass origin outlined above, the lepton number could remain unbroken, making neutrinos Dirac particles. In this scenario, their masses would arise from renormalizable Yukawa interactions, similar to the charged fermions in Eq. (1). The small neutrino masses would then result from extremely tiny Yukawa couplings, further complicating the flavor puzzle.
Flavor model building attempts to explain the non-generic flavor structure of the SM fermions described above. It remains an open question to what extent the flavor hierarchies in the quark sector, the hierarchies in the charged lepton spectrum, and the absence of pronounced hierarchies in the neutrino sector are connected. While this review primarily focuses on the quark sector, we will also discuss the lepton and neutrino sectors where relevant.
@@ 2.2 The New Physics Flavor Puzzle
One practical implication of the SM flavor puzzle, even without knowing its solution, is the existence of approximate accidental symmetries within the SM. The smallness of Yukawa couplings and the CKM alignment leads to approximate flavor and CP conservation, resulting in highly suppressed FCNCs and EDMs. The resulting selection rules, exemplified by the Glashow–Iliopoulos–Maiani (GIM) mechanism [12], have been instrumental in designing precision tests of the SM that are highly sensitive to new physics, which typically breaks these approximate symmetries. So far, these tests have yielded negative results. However, the null findings have profound consequences for the flavor structure of new physics if it exists near the TeV scale—a situation known as the new physics flavor puzzle. That is to say, new flavor parameters of a TeV-scale extension of the SM are also not generic.
@@
Consider a flavor-anarchic scenario in which all dimensionless flavor parameters are of order one, assuming no particular hierarchy or structure in the flavor sector of the new physics. For a tree level and one-loop matching, this predicts $c_{\mathcal{O}}\sim 1$ and $\sim 1/16\pi^{2}$, respectively. Under this assumption, many of the dimension-6 operators contribute significantly to processes involving neutral meson oscillations ($\Delta F=2$), charged lepton flavor violation (cLFV), EDMs, or strangeness-changing decays ($\Delta S=1$). The experimental absence of deviations from SM predictions in such processes imposes stringent constraints on the scale $\Lambda$, often pushing it to values as high as six orders of magnitude above the EW scale. For a concrete illustration, Figure 2 presents bounds on a selected set of operators in the Warsaw basis [13], explicitly showing the flavor indices in the down-quark charged-lepton mass basis. For the dipole operators, we include a one-loop suppression factor expected from the matching, whereas the four-quark operators are considered to be generated at the tree level. For more operator examples, see Ref. [14].
Figure 2: Experimental bounds on a selection of dimension-6 SMEFT operators assuming flavor anarchy. Four-quark operators are considered to be generated at the tree level. Dipole operators contain a loop suppression factor.
@@
The main motivation for expecting new physics at the TeV scale is the Higgs hierarchy problem, which asks why the Higgs boson mass is so much lighter than the Planck scale despite quantum corrections that should drive it higher. Symmetry-based solutions to this problem—such as supersymmetry (SUSY) or composite Higgs models—imply that new physics to stabilize the Higgs mass is expected not far above the electroweak scale. These theories often introduce new particles and interactions with order-one couplings to the Higgs field. Consequently, the flavor structure of such new physics cannot be generic; it must be hierarchical and (or) aligned to avoid excessive violation of the SM’s approximate flavor and CP symmetries. This raises the possibility that the NP and the SM flavor structure might share a common origin.
@@ 3.1.1 Symmetry-Based Models
The Froggatt-Nielsen (FN) mechanism is a seminal example utilizing an abelian horizontal symmetry, typically a $U(1)$ symmetry, to explain flavor hierarchies [15, 16, 17]. In this framework, the SM fermions carry generation-dependent charges under the additional $U(1)$ symmetry. A scalar field, known as the FN (or flavon) field ($\Phi$), acquires a vacuum expectation value (vev) $\langle\Phi\rangle$, thus spontaneously breaking the flavor symmetry. The vev $\langle\Phi\rangle$ is assumed to be smaller than the cutoff scale $\Lambda$ of the effective theory. Yukawa couplings arise from higher-dimensional operators suppressed by powers of the small parameter $\varepsilon=\langle\Phi\rangle/\Lambda\ll 1$. By assigning appropriate $U(1)$ charges, the FN mechanism can qualitatively reproduce the observed mass hierarchies and mixing angles from a single expansion parameter $\varepsilon$. For instance, heavier fermions are assigned lower charges, resulting in less suppression in their Yukawa couplings. In essence, the Yukawa couplings in equation (1) are replaced by the following terms
$\mathcal{L}\supset-\tilde{Y}_{u}^{pr}\varepsilon^{Q_{q_{p}}-Q_{u_{r}}}\,\overline{q}_{p}\widetilde{H}u_{r}-\tilde{Y}_{d}^{pr}\varepsilon^{Q_{q_{p}}-Q_{d_{r}}}\,\overline{q}_{p}Hd_{r}-\tilde{Y}_{e}^{pr}\varepsilon^{Q_{\ell_{p}}-Q_{e_{r}}}\,\overline{\ell}_{p}He_{r}\,,$ (3)
where the $\tilde{Y}_{u,d,e}$ are “proto” Yukawa couplings that can all be of $\mathcal{O}(1)$, and $Q_{q_{p}}$, $Q_{\ell_{p}}$, $Q_{u_{r}}$, $Q_{d_{r}}$, $Q_{e_{r}}$ are the $U(1)$ charges of the quarks and leptons. We assume that the Higgs is uncharged $Q_{H}=0$, and without loss of generality we have set $Q_{\Phi}=1$.
Figure 3: Example of how the Froggatt-Nielsen mechanism might generate a small Yukawa coupling of the down quark. The shown scenario corresponds to the $U(1)$ charges $Q_{q_{1}}=+3$, $Q_{d_{1}}=-3$, $Q_{H}=0$, and $Q_{\Phi}=+1$. Left: Chain of vectorlike fermions $\psi_{i}$ that connect the down quarks to the Higgs boson. Right: the corresponding higher dimensional operator. Figure adapted from Reference [7], with permission from the author.
The most straightforward UV completion of the higher dimensional operators involves chains of vector-like fermions that communicate the symmetry breaking to the SM fermions. In such a case, the scale $\Lambda$ corresponds to the masses of the vector-like fermions, and the proto Yukawas are combinations of their couplings. A typical example is shown in Figure 3. The FN setup contains a number of free $\mathcal{O}(1)$ parameters sufficient to reproduce the observed fermion masses precisely. If the $U(1)$ symmetry is gauged, the models contain a $Z^{\prime}$ gauge boson with a mass of order $\langle\Phi\rangle$. If the $U(1)$ is global, one instead expects a very light pseudo-Nambu-Goldstone boson. These particles generically have flavor-changing couplings and can be effectively searched for with flavor probes. Flavor models commonly feature a decoupling limit: by sending both $\langle\Phi\rangle$ and $\Lambda$ to high energies while keeping $\varepsilon=\langle\Phi\rangle/\Lambda$ constant, these models can address the SM flavor puzzle without introducing excessive FCNCs at low energies. Current experimental constraints impose a lower bound on the scale of flavor dynamics, which are, however, model-dependent. In the simplest cases with gauged $U(1)$ symmetries, the $Z^{\prime}$ can contribute at the tree level to kaon mixing, and bounds can be as stringent as $\langle\Phi\rangle\gtrsim 10^{4}$ TeV. In the case of a global $U(1)$, one often finds even stronger bounds on the new physics scale from flavor-violating decays of kaons and muons into the light pseudo-Nambu-Goldstone boson, see section 4.5 for more details. Interestingly, for a discrete FN, the breaking scale can be as low as $\langle\Phi\rangle\sim$ TeV; see [18]. Future experiments could potentially reveal evidence for low-scale flavor models.
@@
Non-abelian flavor symmetries, such as $SU(3)$ or $SU(2)$, organize the fermion generations into non-trivial group representations. By assigning fermions to multiplets, these symmetries constrain the structure of the Yukawa couplings. The observed flavor structure typically arises through sequential and hierarchical symmetry breaking. Non-abelian symmetries often lead to more restrictive textures than FN models. However, challenges to reproducing realistic parameters in these models include the complexity of the symmetry-breaking sector. Models based on $SU(3)$ flavor symmetry have been proposed to generate realistic mass hierarchies; see e.g. [19, 20, 21, 22]. However, the sizable Yukawa coupling of the top quark ($y_{t}\sim 1$) breaks the $SU(3)$ symmetry, making $SU(2)$ (or better $U(2)$) groups more attractive [23, 24, 25, 26, 27, 28, 29]. In this construction, the first two families form a $U(2)$ doublet, while the third family is a singlet. In the symmetric limit, only the third generation acquires mass, providing a decent starting point. Symmetry breaking typically occurs in two stages: first, the $U(2)\equiv SU(2)\times U(1)$ symmetry is broken to an intermediate $U(1)$, generating a small mass for the second generation; subsequently, the $U(1)$ symmetry is broken, yielding an even smaller mass for the first generation. For recent developments, see section 4.1.
@@
Radiative models propose that lighter SM fermion masses arise from loop corrections rather than at the tree level [30]. In these scenarios, mass hierarchies arise from the suppression associated with loop factors, which intriguingly coincide with the two orders of magnitude separating the masses of adjacent generations. The absence of tree level Yukawa couplings for the light generations can, for example, be imposed by using flavor symmetries. The breaking of a flavor symmetry is then communicated from a separate sector of the model to the SM fermions via loops. Additional fields and interactions are introduced to facilitate the necessary loop diagrams. In contrast to FN models, their masses can be at the same scale as flavor breaking, as the role of $\varepsilon$ is effectively replaced by a loop factor. The wide range of possible quantum numbers and interactions for the new fields gives rise to a plethora of models, both supersymmetric [31, 32, 33, 34] and non-supersymmetric [35, 36, 37, 38].
@@ 3.1.2 Geometry-Based Models
A concrete example in which this mechanism can be at work is the Randall-Sundrum (RS) model [40] that introduces a warped extra dimension with a non-factorizable geometry, where the metric depends exponentially on the extra-dimensional coordinate. The extra dimension is bounded by an infrared (IR) brane and an ultraviolet (UV) brane. The warping leads to exponential factors that naturally generate hierarchies. The Higgs naturalness problem is addressed if the Higgs resides on the IR brane. Fermions that propagate in the bulk have profiles that interpolate between the UV and IR branes and can be localized at different positions along the extra dimension, resulting in hierarchical Yukawa couplings with the Higgs. In such a case, the effective Yukawa couplings are given by
$\mathcal{L}\supset-\hat{Y}_{u}^{pr}f_{q_{p}}f_{u_{r}}\,\overline{q}_{p}\widetilde{H}u_{r}-\hat{Y}_{d}^{pr}f_{q_{p}}f_{d_{r}}\,\overline{q}_{p}Hd_{r}-\hat{Y}_{e}^{pr}f_{\ell_{p}}f_{e_{r}}\,\overline{\ell}_{p}He_{r}\,,$ (4)
where the $\hat{Y}_{u,d,e}$ are the 5-dimensional Yukawa couplings that can be $\mathcal{O}(1)$, and the $f_{\Psi}$ are the wavefunctions evaluated on the IR brane. One finds
$f_{\Psi}\simeq\begin{cases}\sqrt{1-2c_{\Psi}}~{},\quad&\text{for}~{}~{}c_{\psi}<1/2~{},\\ \sqrt{2c_{\Psi}-1}~{}e^{-(c_{\Psi}-1/2)L}~{},\quad&\text{for}~{}~{}c_{\psi}>1/2~{},\\ \end{cases}$ (5)
where $L$ is the size of the extra dimension and the $c_{\Psi}$ are called bulk masses that are free parameters. Heavier fermions, such as the top quark, are more localized towards the IR brane ($c_{\Psi}<1/2$), resulting in $\mathcal{O}(1)$ Yukawa couplings, while light fermions are localized towards the UV brane ($c_{\Psi}>1/2$) resulting in exponentially small Yukawas [41, 42, 43]. A typical example is shown in Figure 4.
Figure 4: Illustration of the up quark and top quark extra-dimensional wave function profiles in a Randall-Sundrum model. The up quark is localized towards the UV brane, while the top towards the IR brane. The chosen bulk mass parameters are $c_{u}=0.7$ and $c_{t}=0.2$. The Higgs is shown localized on the IR brane. The size of the extra dimension is set to $L\simeq 37$ in Planck units, such that $e^{L}\simeq M_{\text{Pl}}/(1\,\text{TeV})$.
The flavor-dependent, extra-dimensional profiles of the SM fermions imply a rich flavor phenomenology [44, 45]. In generic setups with warped extra dimensions, the constraints from meson mixing are very stringent and push the new physics scale to $\mathcal{O}(10)$ TeV, thus creating an obstacle for a fully natural solution of the electroweak hierarchy problem. Extra dimensional models can be protected from large FCNCs by using bulk and brane flavor symmetries [46]. However, in those models, the flavor hierarchies are no longer generated by the attractive mechanism of wave function overlap.
Partial compositeness proposes that the SM fermions are mixtures of elementary and composite states arising from a strongly coupled sector not far above the electroweak scale. Models of partial compositeness were originally introduced to tame FCNCs in technicolor models [47]. The light fermions are mostly elementary, while the heavy top quark is mostly composite. Per the AdS/CFT correspondence, models of partial compositeness map onto models with warped extra dimensions, with the amount of compositeness given by the localization of the fermions along the extra dimension.